# MSL: Not All Tokens Are What You Need for Tuning LLM as a Recommender

### v1

@@ Abstract.
Sequential Recommendation, Large Language Model, Loss Function
ccs: Information systems Recommender systems
@@ 1. Introduction
Figure 1. Language modeling loss can erroneously treat all fictional items as negative items. However, some of these items may exhibit semantic similarities to positive items (e.g., ”Iron Man: AI Rebellion”). Additionally, certain items might are fictional simply because they have not yet been released (e.g., ”Iron Man 4”). These cases represent incorrect negative signals.
@@ 3. Analyses on Language Modeling Loss
(3) $\mathcal{L}_{LML}^{1}(x,y^{p};\theta)$ $=\sum_{t=1}^{|y^{p}|}\underbrace{-\log\frac{\sum_{z\in\mathcal{Z}_{valid}(y^{p}_{<t})}\exp(f_{\theta}(z)))}{\sum_{z\in\mathcal{Z}}\exp(f_{\theta}(v)))}}_{\text{Lifting valid items over invalid items}}$
(4) $\mathcal{L}_{LML}^{2}(x,y^{p};\theta)$ $=\sum_{t=1}^{|y^{p}|}\underbrace{-\log\frac{\exp(f_{\theta}(y_{t}^{p})))}{\sum_{z\in\mathcal{Z}_{valid}(y^{p}_{<t})}\exp(f_{\theta}(z)))}}_{\text{Lifting the positive items over negative items}}$
@@ 4.1. Masked Softmax Loss
(5) $\mathcal{L}_{MSL}(x,y^{p};\theta)=\sum_{t=1}^{|y^{p}|}-\log\frac{\exp(f_{\theta}(y_{t}^{p})))}{\sum_{z\in\mathcal{Z}_{valid}}\exp(f_{\theta}(z)))}$
@@
$P_{\theta}^{valid}(y^{v}_{t}|x,y^{v}_{<t})=\frac{\exp(f_{\theta}(y_{t}^{v})))}{\sum_{z\in\mathcal{Z}_{valid}(y^{v}_{<t})}\exp{(f_{\theta}(z))}}$
@@ 4.2.2. The Introduction of Temperature
(9) $\mathcal{L}_{MSL}(x,y^{p};\theta)=\sum_{t=1}^{|y^{p}|}-\tau\log\frac{\exp(f_{\theta}(y_{t}^{p})/\tau))}{\sum_{z\in\mathcal{Z}_{valid}}\exp(f_{\theta}(z)/\tau))}$
@@
(10) $w(y_{t}^{p})=1-P^{valid}_{\theta}(y_{t}^{p})=1-\frac{\exp(f_{\theta}(y_{t}^{p})/\tau))}{\sum_{z\in\mathcal{Z}_{valid}}\exp{(f_{\theta}(z)/\tau)}}$
@@ Acknowledgements.
(13) $\mathcal{L}_{MSL}(x,y^{p};\theta)$ $=\sum_{t=1}^{|y^{p}|}-\log\frac{\exp(f_{\theta}(y_{t}^{p})))}{\sum_{z\in\mathcal{Z}_{valid}}\exp(f_{\theta}(z)))}$
@@
(15) $\mathcal{L}_{MSL}(x,y^{p};\theta)$ $\geq-\log\frac{1}{\sum_{v\in\mathcal{V}}\mathbb{I}(\log P_{\theta}^{valid}\left(y^{p}\mid x\right)\geq\log P_{\theta}^{valid}\left(y^{v}\mid x\right))}$
@@
Since the denominator of the softmax function in each term of $\mathcal{L}_{{MSL}}$ is smaller than $\mathcal{L}_{LML}$, it follows that $\mathcal{L}_{MSL}<\mathcal{L}_{LML}$, then

### v4

@@ Abstract.
Sequential Recommendation; Large Language Model; Loss Function
ccs: Information systems Recommender systemsjournalyear: 2025copyright: acmlicensedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730041isbn: 979-8-4007-1592-1/2025/07
@@ 1. Introduction
Figure 1. Language modeling loss can erroneously treat all fictional items as negative items. However, some of these items may exhibit semantic similarities to positive items (e.g., ”Iron Man: AI Rebellion”). Additionally, certain items might be fictional simply because they have not yet been released (e.g., ”Iron Man 4”). These cases represent incorrect negative signals.
@@ 3. Analyses on Language Modeling Loss
(3) $\mathcal{L}_{LML}^{1}(x,y^{p};\theta)$ $=\sum_{t=1}^{|y^{p}|}\underbrace{-\log\frac{\sum_{z\in\mathcal{Z}_{valid}(y^{p}_{<t})}\exp(f_{\theta}(z))}{\sum_{z\in\mathcal{Z}}\exp(f_{\theta}(z))}}_{\text{Lifting valid items over invalid items}}$
(4) $\mathcal{L}_{LML}^{2}(x,y^{p};\theta)$ $=\sum_{t=1}^{|y^{p}|}\underbrace{-\log\frac{\exp(f_{\theta}(y_{t}^{p}))}{\sum_{z\in\mathcal{Z}_{valid}(y^{p}_{<t})}\exp(f_{\theta}(z))}}_{\text{Lifting the positive items over negative items}}$
@@ 4.1. Masked Softmax Loss
(5) $\mathcal{L}_{MSL}(x,y^{p};\theta)=\sum_{t=1}^{|y^{p}|}-\log\frac{\exp(f_{\theta}(y_{t}^{p}))}{\sum_{z\in\mathcal{Z}_{valid}}\exp(f_{\theta}(z))}$
@@
$P_{\theta}^{valid}(y^{v}_{t}|x,y^{v}_{<t})=\frac{\exp(f_{\theta}(y_{t}^{v}))}{\sum_{z\in\mathcal{Z}_{valid}(y^{v}_{<t})}\exp{(f_{\theta}(z))}}$
@@ 4.2.2. The Introduction of Temperature
(9) $\mathcal{L}_{MSL}(x,y^{p};\theta)=\sum_{t=1}^{|y^{p}|}-\tau\log\frac{\exp(f_{\theta}(y_{t}^{p})/\tau)}{\sum_{z\in\mathcal{Z}_{valid}}\exp(f_{\theta}(z)/\tau)}$
@@
(10) $w(y_{t}^{p})=1-P^{valid}_{\theta}(y_{t}^{p})=1-\frac{\exp(f_{\theta}(y_{t}^{p})/\tau)}{\sum_{z\in\mathcal{Z}_{valid}}\exp{(f_{\theta}(z)/\tau)}}$
@@ Acknowledgements.
(13) $\mathcal{L}_{MSL}(x,y^{p};\theta)$ $=\sum_{t=1}^{|y^{p}|}-\log\frac{\exp(f_{\theta}(y_{t}^{p}))}{\sum_{z\in\mathcal{Z}_{valid}}\exp(f_{\theta}(z))}$
@@
(15) $\mathcal{L}_{MSL}(x,y^{p};\theta)$ $\geq-\log\frac{1}{\sum_{v\in\mathcal{V}}\mathbb{I}(\log P_{\theta}^{valid}\left(y^{v}\mid x\right)\geq\log P_{\theta}^{valid}\left(y^{p}\mid x\right))}$
@@
Since the denominator of the softmax function in each term of $\mathcal{L}_{{MSL}}$ is smaller than $\mathcal{L}_{LML}$, it follows that $\mathcal{L}_{MSL}\leq\mathcal{L}_{LML}$, then